40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Apr 18 22:35:23 2019, gautam, Update, SUS, ETMY actuator diagnosis  
    Reply  Fri Apr 19 11:36:23 2019, gautam, Update, SUS, No consistent solution for output matrix 
       Reply  Fri Apr 19 15:13:38 2019, rana, Update, SUS, No consistent solution for output matrix 
          Reply  Fri Apr 19 16:19:42 2019, gautam, Update, SUS, Actuation matrix still not orthogonal 
             Reply  Fri Apr 19 19:22:15 2019, rana, Update, SUS, Actuation matrix still not orthogonal 
    Reply  Mon Apr 22 22:43:15 2019, gautam, Update, SUS, ETMY sensor diagnosis  ETMY_sensorSpectra_consolidated.pdf
       Reply  Thu Apr 25 00:30:45 2019, gautam, Update, SUS, ETMY BR mode  ETMY_sensorSpectra_BRmode.pdfETMY_sensorSpectra_BRmode.pdfETMX_sensorSpectra_BRmode.pdfIMG_5993.JPG
Message ID: 14569     Entry time: Thu Apr 25 00:30:45 2019     In reply to: 14562
Author: gautam 
Type: Update 
Category: SUS 
Subject: ETMY BR mode  

We briefly talked about the bounce and roll modes of the SOS optic at the meeting today. 

Attachment #1: BR modes for ETMY from my free-swinging run on 17 April. The LL coil has a very different behavior from the others.

Attachment #2: BR modes for ETMY from my free-swinging run on 18 April, which had a macroscopically different bias voltage for the PIT/YAW sliders. Here too, the LL coil has a very different behavior from the others.

Attachment #3: BR modes for ETMX from my free-swinging run on 27 Feb. There are many peaks in addition to the prominent ones visible here, compared to ITMY. The OSEM PD noise floor for UR and SIDE is mysteriously x2 lower than for the other 3 OSEMs???

In all three cases, a bounce mode around 16.4 Hz and a roll mode around 24.0 Hz are visible. The ratio between these is not sqrt(2), but is ~1.46, which is ~3% larger. But when I look at the database, I see that in the past, the bounce and roll modes were in fact at close to these frequencies.

In conclusion:

  1. the evidence thus far says that ETMY has 5 resonant modes in the free-swinging data between 0.5 Hz and 25 Hz.
  2. Either two modes are exactly degenerate, or there is a constraint in the system which removes 1 degree of freedom.
  3. How likely is the latter? As any mechanical constraint that removes one degree of freedom would presumably also damp the Qs of the other modes more than what we are seeing.
  4. Can some large piece of debris on the barrel change the PIT/YAW eigenvectors such that the eigenvalues became exactly degenerate?
  5. Furthermore, the AC actuation vectors for PIT and YAW are not close to orthogonal, but are rotated ~45 degrees relative to each other.

Because of my negligence and rushing the closeout procedure, I don't have a great close-out picture of the magnet positions in the face OSEMs, the best I can find is Attachment #4. We tried to replicate the OSEM arrangement (orientation of leads from the OSEM body) from July 2018 as closely as possible.

I will investigate the side coil actuation strength tomorrow, but if anyone can think of more in-air tests we should do, please post your thoughts/poetry here.

Attachment 1: ETMY_sensorSpectra_BRmode.pdf  130 kB  Uploaded Thu Apr 25 01:34:38 2019  | Hide | Hide all
ETMY_sensorSpectra_BRmode.pdf
Attachment 2: ETMY_sensorSpectra_BRmode.pdf  153 kB  Uploaded Thu Apr 25 01:34:52 2019  | Hide | Hide all
ETMY_sensorSpectra_BRmode.pdf
Attachment 3: ETMX_sensorSpectra_BRmode.pdf  176 kB  Uploaded Thu Apr 25 01:35:05 2019  | Hide | Hide all
ETMX_sensorSpectra_BRmode.pdf
Attachment 4: IMG_5993.JPG  950 kB  Uploaded Thu Apr 25 01:51:25 2019  | Hide | Hide all
IMG_5993.JPG
ELOG V3.1.3-